• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Puckett

    Log entry time 08:00:51 on April18,2007

    Entry number 199474

    keyword=Owl Shift Summary

    Started in same running conditions as yesterday. Plan according to previous shift is to keep running this way all night.

    Runs (Run #, Start-Stop, N, Replay Done?)
    2603, 23:57-00:24, 1.2M, Y
    2604, 00:25-00:51, 1.2M, Y
    2605, 00:51-01:15, 1.2M, Y
    2606, 01:16-01:38, 557K, Y

    Beam down due to RF ON problems with a cavity in the NL, lasts about 23 minutes

    2607, 01:51-02:17, 1.2M, Y
    2608, 02:18-02:45, 1.2M, Y
    2609, 02:45-03:48, 0.9M, Y

    CODA problem: communication with ROCs and TS were still okay, but no events were coming in, even though the beam was on. It is likely we lost a little over 1/2 hour as a result. We killed CODA and restarted and now we are back up and running. Will keep an eye on it.

    2610, 03:51-04:17, 1.2M, Y
    2611, 04:17-04:43, 1.2M, Y
    2612, 04:43-05:10, 1.2M, Y
    2613, 05:11-05:40, 1.2M, CODA crash upon reaching event limit, TS11 stopped responding, no end-of-run file!
    2614, 05:47-06:14, 1.2M, Y
    2615, 06:15-06:44, 1.2M, Y
    2616, 06:48-07:12, 1.2M, Y
    2617, 07:13-07:40, 1.2M, Y
    2618, 07:40-??